OpenStudy (anonymous):

When a ball is thrown upward, the formula h=-16t^2+64t gives its height h (in feet) above the ground after t seconds. How many seconds does it take for the ball to strike the ground?

OpenStudy (anonymous):

solve for t wen h = 0 :P

OpenStudy (anonymous):

i dont know how to do that :(

OpenStudy (anonymous):

-16t(t-4) = 0

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

solve that equation using the zero product property...

OpenStudy (anonymous):

16t=0 and t=4

OpenStudy (anonymous):

because when the ball hits the ground height=0

OpenStudy (anonymous):

im lost :(

OpenStudy (anonymous):

-16t(t-4) = 0 wen t = 0 r t-4 = 0 u get t= 0 r 4 ok?

OpenStudy (anonymous):

lol im kind of slow , can u write out the equation so i can solve it?

OpenStudy (anonymous):

ok, when the ball hits the ground, the height is 0 because the ball is on the ground, so make the equation equal to zero and the factor it out as shown above. You will have two solutions since you have a t^2 which will be 16t=0 and t=4

OpenStudy (anonymous):

\[16t ^{2}+64t= 16t(t+4)=0\] Now your values will be t=0 and when t=-4 but since time can't be negative, it's t=0 only.

OpenStudy (anonymous):

actually its 16t(t-4) and then you have t=4 and t=0 apologies..

OpenStudy (anonymous):

i still dont get it but i'll figure something out, thanks .

OpenStudy (anonymous):

cmon.......u need time wen ball hits the ground r8

OpenStudy (anonymous):

Ok assume you throw a ball upwards, it will get to a maximum height and then it will come down again right? This is how it's path would look like Forming a parabola. So when the ball comes back down, and hits the ground the height of the ball will be zero since it's on the ground. So to find the time it takes to hit the ground, we find the time when the height is equal to zero. So you just put h as o and solve for t. Factoring out -16t and then you would have -16t(t-4) = 0 Equations of such form give you two answers, 16t=0 which is just t=0 and t-4=0 which is t=4 I hope that helps.

OpenStudy (anonymous):

i get it now thanks ! it was much easier when i put it into my graphing calculator :) @beeqay